Find the work done pulling a sled 75๐‘š with a 70๐‘ force if the handle makes an angle of 36ยฐ with the ground.ย 


1
Expert's answer
2021-04-29T15:26:01-0400

Ans:

Given "F=70N" "S=75m" Angle between them is "36\degree"

We know that

Work done by this force will be

"W=F\times S\times Cos36\degree"

"=70\times 75\times0.8090"


"=4247.25J"